Given the computed mean value, µ = 85 and the standard deviation, σ = 5, convert the following raw scores to its z-scores:

a.) x = 92


1
Expert's answer
2022-02-28T12:25:32-0500
"z=\dfrac{x-\mu}{\sigma}=\dfrac{92-85}{5}=1.4"